# Fast adaptive discontinuous basis sets for electronic structure

### v1

@@ Abstract
We develop a discontinuous Galerkin (DG) framework for automatically constructing adaptive basis sets for electronic structure calculations. By allowing basis functions to be discontinuous across element interfaces, our approach supports flexible combinations of atom-centered and polynomial basis sets, maintains favorable numerical conditioning, and induces structured sparsity of the one- and two-electron integrals, which we compute using specialised numerical integration strategies. We also introduce multigrid-preconditioned Poisson solvers that enable fast algorithms for both Hartree-Fock (HF) and density functional theory (DFT) calculations within our DG basis sets. Moreover, these basis sets naturally support adaptive multigrid preconditioning for the linear eigensolvers employed within the self-consistent field iteration for HF and DFT. Numerical experiments for HF and DFT demonstrate that our approach achieves chemical accuracy with modest basis sizes that compare favorably to the sizes of ordinary GTO basis sets achieving similar accuracy, while offering additional structured sparsity and improved computational scalability in the size-extensive limit. The framework thus provides a flexible route toward the construction of systematically improvable and structured adaptive basis sets for electronic structure theory.
@@ 1 Introduction
In this work, we introduce an alternative DG framework for discretising the electronic structure Hamiltonian. We construct basis sets adapted to the problem geometry that are built from primitive functions consisting possibly of both GTOs and polynomials, with supports suitably restricted to individual elements. A straightforward adaptive filtering procedure is applied to ensure orthogonality and to control the size of the resulting basis sets.
@@ 3 Discontinuous Galerkin framework
Now we provide the details of the Discontinuous Galerkin (DG) framework. In particular, we focus on the Symmetric Interior Penalty (SIP) method for discretising the Laplacian operator using discontinuous basis functions. Additionally, we discuss strategies for selecting the penalty parameter to ensure the stability and accuracy of the discretisation.
@@ 8.2 Adaptive basis truncation
For each of the molecules that we consider, we set up a experiment where a provisional basis is constructed using cc-pVQZ basis functions for each atom, with supports suitably restricted to individual elements. Then we perform a HF or DFT calculation using this full DG basis with no filtration, furnishing a reference energy $E_{0}$.
@@ 8.3 $\boldsymbol{\mathrm{H}_{2}}$
$c$-pVDZ 30 -1.090205 0.001423 16 -1.089673 0.001920
$c$-pVTZ 88 -1.091334 0.000258 46 -1.091166 0.000428
$c$-pVQZ 100 -1.091580 0.000059 108 -1.091481 0.000113
@@
$c$-pVDZ 33 -1.106409 0.001492 16 -1.106249 0.001651
$c$-pVTZ 88 -1.107545 0.000356 46 -1.107502 0.000399
$c$-pVQZ 100 -1.107804 0.000097 108 -1.107791 0.000110
@@
For a given level of GTO approximation, lower energies are obtained using the DG framework (where the GTOs are used to create a provisional DG basis), compared to the PySCF calculation (where the GTOs are used directly as continuous basis functions). This improvement can be attributed to the greater expressiveness of the discontinuous basis, as GTOs are duplicated and restricted in our framework, such that each provisional basis function has support only on a single element. The greater flexibility of the DG basis permits more effective representation of interstitial regions between atomic nuclei.
@@ 8.4 $\boldsymbol{\mathrm{LiH}}$
Again we observe that in the DZ and TZ cases, the induced DG basis is noticeably larger than the ordinary GTO basis, but this trend reverses for the larger QZ basis. However, in contrast with the previous example, we do not see any clear gains in accuracy compared to PySCF from simply cutting GTOs on each element. One possible reason is that the $\mathrm{LiH}$ bond is ionic, and the electron density is localised to the atomic nuclei.
@@
$c$-pVDZ 73 -7.985820 0.001517 43 -7.985790 0.001547
$c$-pVTZ 150 -7.986966 0.000371 94 -7.986964 0.000373
$c$-pVQZ 180 -7.987182 0.000145 198 -7.987228 0.000109
@@
$c$-pVDZ 73 -7.918130 0.001465 43 -7.918077 0.001519
$c$-pVTZ 150 -7.919294 0.000302 94 -7.919217 0.000378
$c$-pVQZ 180 -7.919462 0.000134 198 -7.919493 0.000103
@@
We also examine the energies of the first three virtual orbitals obtained from HF, depicted in Figure 12. For this experiment we have used the cc-pVQZ basis, both directly and as a means to construct a DG basis. As a reference, we compare against the virtuals obtained using the aug-cc-pVQZ GTO basis in PySCF.
@@ 8.5 $\boldsymbol{\mathrm{H}_{2}\mathrm{O}}$
We do observe that the DG basis sets yield more accurate energies than the corresponding ordinary GTO basis sets, consistent with previous examples exhibiting covalent bonding.
@@
$c$-pVDZ 105 -76.032872 0.034171 51 -76.030384 0.036659
$c$-pVTZ 220 -76.059495 0.007548 114 -76.057214 0.009829
$c$-pVQZ 294 -76.066381 0.000662 255 -76.064823 0.002220
@@
$c$-pVDZ 105 -75.871874 0.040787 51 -75.867762 0.044899
$c$-pVTZ 220 -75.902357 0.010304 114 -75.910937 0.013143
$c$-pVQZ 294 -75.910937 0.001724 255 -75.909232 0.003429
@@ 8.6 $\boldsymbol{\mathrm{C}_{6}\mathrm{H}_{6}}$
The results are similar to the previous cases exhibiting covalent bonding. For lower levels of GTO approximation, the induced DG basis is smaller than the ordinary GTO basis set, but the DG basis is actually smaller in the QZ case. Likewise, the DG basis sets for a given level of GTO approximation yield more accurate energies.
@@
$c$-pVDZ 476 -230.136403 0.065688 258 -230.135675 0.066417
$c$-pVTZ 966 -230.188860 0.013232 552 -230.184766 0.017326
$c$-pVQZ 1140 -230.200599 0.001492 1188 -230.198799 0.003292
@@ 9 Conclusion
Interestingly, we find that combining polynomial functions with GTOs in the provisional basis offers no significant accuracy gain over using GTOs alone. Nonetheless, polynomials offer a systematic path toward achieving a complete basis, which GTOs alone do not provide. An important future direction is to explore the question of systematic convergence for post-HF methods, where basis set incompleteness errors are more pronounced.

### v2

@@ Abstract
We develop a discontinuous Galerkin (DG) framework for automatically constructing adaptive basis sets for electronic structure calculations. By allowing basis functions to be discontinuous across element interfaces, our approach supports flexible combinations of atom-centered and polynomial basis sets, maintains favorable numerical conditioning, and induces structured sparsity of the one- and two-electron integrals, which we compute using specialised numerical integration strategies. In addition, we introduce a simple post-processing procedure to obtain continuous solutions if desired. We also introduce multigrid-preconditioned Poisson solvers that enable fast algorithms for both Hartree-Fock (HF) and density functional theory (DFT) calculations within our DG basis sets. Moreover, these basis sets naturally support adaptive multigrid preconditioning for the linear eigensolvers employed within the self-consistent field iteration for HF and DFT. Numerical experiments for HF and DFT demonstrate that our approach achieves chemical accuracy with modest basis sizes that compare favorably to the sizes of ordinary GTO basis sets achieving similar accuracy, while offering additional structured sparsity and improved computational scalability in the size-extensive limit. The framework thus provides a flexible route toward the construction of systematically improvable and structured adaptive basis sets for electronic structure theory.
@@ 1 Introduction
In this work, we introduce an alternative DG framework for discretising the electronic structure Hamiltonian. We construct basis sets adapted to the problem geometry that are built from primitive functions consisting possibly of both GTOs and polynomials, with supports suitably restricted to individual elements. A straightforward adaptive filtering procedure is applied to ensure orthogonality and to control the size of the resulting basis sets. Finally, to obtain fully continuous solutions using the discontinuous basis sets, we introduce an optional projection onto the continuous subspace embedded in the overarching discontinuous function space.
@@ 3 Discontinuous Galerkin framework
Now we provide the details of the Discontinuous Galerkin (DG) framework. In particular, we focus on the Symmetric Interior Penalty (SIP) method for discretising the Laplacian operator using discontinuous basis functions. Additionally, we discuss strategies for selecting the penalty parameter to ensure the stability and accuracy of the discretisation. Finally, we describe a simple post-processing procedure to project any discontinuous function discretised in our DG framework to the nearest continuous function.
@@ 3.5 Penalty parameter
3.6 Continuous projection
Consider isolating the penalty matrix from the discrete Laplacian:
$P_{ij}=\int_{\Gamma}[[\phi_{i}]]\cdot[[\phi_{j}]]\penalty 10000\ ds,$ (23)
and note that the basis coefficients of any continuous function in $V_{\mathrm{DG}}$ lie in the null space of $P$. As such, given any function $v\in V_{\mathrm{DG}}$, we can obtain the continuous part via orthogonal projection onto the null space of $P$.
As $P$ is a symmetric positive semi-definite matrix, the projection $\Pi$ onto the null space of $P$ can be computed as
$\Pi=I-P^{+}P=I-\lim_{\varepsilon\rightarrow 0}(P+\varepsilon I)^{-1}P=\lim_{\varepsilon\rightarrow 0}\varepsilon(P+\varepsilon I)^{-1}.$ (24)
In practice, we pick a small value of $\varepsilon=10^{-3}$ and evaluate the action of the limiting expression
$\Pi_{\varepsilon}:=\varepsilon(P+\varepsilon I)^{-1}$ (25)
with a linear solve. To perform the linear solve, we apply the conjugate gradient algorithm preconditioned with an incomplete Cholesky factorisation with zero additional fill. We find that this simple approach is sufficient for our numerical tests as it is only applied as a post-processing step and does not present as a computational bottleneck. Moreover, by contrast, incomplete Cholesky for our Laplacian solves performs poorly.
To understand how $\varepsilon$ controls the failure of continuity, let $c$ denote the vector of coefficients of an $L^{2}$-normalized function in our DG basis, so that $\|c\|_{2}=1$, and let $c_{\varepsilon}:=\Pi_{\varepsilon}c$ denote the coefficients after projection. Then the ‘discontinuity seminorm’ after projection can be bounded as
$\sqrt{c_{\varepsilon}^{\top}Pc_{\varepsilon}}\leq\varepsilon/2.$ (26)
Indeed, $c_{\varepsilon}^{\top}Pc_{\varepsilon}=c^{\top}f_{\varepsilon}(P)c$, where $f_{\varepsilon}(P)$ denotes the matrix function induced by the univariate function $f_{\varepsilon}(x)=\varepsilon^{2}x/(x+\varepsilon^{2})$, which satisfies $f_{\varepsilon}(x)\leq\varepsilon/4$ for all $x\geq 0$. The discontinuity seminorm after projection may in fact be significantly smaller, since $c$ typically represents a function that is already nearly continuous.
In practice, we find that post-processing our solutions using such an approximate projection only negligibly affects our estimates of Hartree-Fock and DFT energies, but we still include the step due to the relatively insignificant cost and the reassurance of approximately restoring variational energy guarantees that are automatic for continuous Galerkin approaches.
Indeed, for the purpose of estimating energies, note that the action of the discrete DG Laplacian matrix coincides with the that of the continuous part
$T_{ij}^{\mathrm{cont}}=\sum_{m=1}^{M}\int_{K_{m}}\nabla\phi_{i}(\mathbf{r})\cdot\phi_{j}(\mathbf{r})\penalty 10000\ d\mathbf{r}.$ (27)
on the null space of $\Pi$. After the approximate projection is applied to define new normalized basis coefficients ${c}_{\varepsilon}:=\Pi_{\varepsilon}c$, we compute the kinetic energy using $T^{\mathrm{cont}}$ in place of $T$. In our implementation, we verify that
$\frac{{c}_{\varepsilon}^{\top}(T-T^{\mathrm{cont}}){c}_{\varepsilon}}{\|{c}_{\varepsilon}\|^{2}}<10^{-8}$ (28)
to ensure that the use of $T^{\mathrm{cont}}$ does not affect the calculation of energies.
@@ 8.2 Adaptive basis truncation
For each of the molecules that we consider, we set up a experiment where a provisional basis is constructed using cc-pVQZ basis functions for each atom, with supports suitably restricted to individual elements. Then we perform a HF or DFT calculation using this full DG basis with no filtration, furnishing a reference energy $E_{0}$. Finally, we apply the procedure in Section 3.6 to obtain an approximately continuous wavefunction which is used to evaluate the energy.
We note in all our numerical experiments that with our choice of penalty parameter following the discussion in Section 3.5, application of the continuous projection does not alter the obtained wavefunctions by more than $10^{-5}$ in $L^{2}$ norm in practice. As such we only report the energies associated with the continuous projected wavefunctions.
@@ 8.3 $\boldsymbol{\mathrm{H}_{2}}$
$c$-pVDZ 30 -1.090161 0.001467 16 -1.089673 0.001920
$c$-pVTZ 88 -1.091264 0.000374 46 -1.091166 0.000428
$c$-pVQZ 100 -1.091529 0.000109 108 -1.091481 0.000113
@@
$c$-pVDZ 33 -1.106407 0.001494 16 -1.106249 0.001651
$c$-pVTZ 88 -1.107538 0.000363 46 -1.107502 0.000399
$c$-pVQZ 100 -1.107799 0.000102 108 -1.107791 0.000110
@@
For a given level of GTO approximation, slightly lower energies are obtained using the DG framework (where the GTOs are used to create a provisional DG basis), compared to the PySCF calculation (where the GTOs are used directly as continuous basis functions). This improvement can be attributed to the greater expressiveness of the discontinuous basis, as GTOs are duplicated and restricted in our framework, such that each provisional basis function has support only on a single element. The greater flexibility of the DG basis permits more effective representation of interstitial regions between atomic nuclei.
@@
8.3.1 Mesh dependence
We investigate the dependence of the computed energy on the underlying DG mesh by performing a mesh-sensitivity experiment on the $\mathrm{H}_{2}$ molecule. The two hydrogen atoms are centered one unit apart $(0,0,0)$ and $(1,0,0)$.
The computational domain is partitioned into two DG elements (one associated with each atom), separated by an inter-element boundary located at position $x=h$. By varying $h\in[0,1]$, we slide the interface between the two atoms while keeping the nuclear positions fixed. For each choice of $h$, we construct the corresponding DG mesh and compute the HF and DFT energies using the cc-pVQZ basis as the provisional basis set.
A schematic of this setup and the resulting energies as a function of the interface position $h$ are reported in Figure 10. We observe that the HF and DFT energies to vary only very slightly, on the order of $10^{-5}$, as $h$ is varied. This observation extends all the way to $h=0$ (or, by symmetry, $h=1$), where the interface sits on one of the atoms. As such, we do not expect our calculated energies to depend heavily on the initial mesh construction.
$h\in[0,1]$h=0h=0h=1h=1HH
HHh=0.25h=0.25
00.250.250.50.51.09158-1.091581.09156-1.091561.09154-1.091541.09152-1.091521.0915-1.09151.09148-1.091481.09146-1.09146hhEnergy (HF)
00.250.250.50.51.10784-1.107841.10782-1.107821.1078-1.10781.10778-1.107781.10776-1.10776hhEnergy (DFT)
Figure 10: Top: schematic for the mesh sensitivity test, where the inter-element interface is positioned at $x=h\in[0,1]$. Bottom: HF and DFT energies as a function of $h\in[0,1/2]$, obtained using cc-pVQZ bas the provisional basis set.
@@ 8.4 $\boldsymbol{\mathrm{LiH}}$
Again we observe that in the DZ and TZ cases, the induced DG basis is noticeably larger than the ordinary GTO basis, but this trend reverses for the larger QZ basis. We also examine the energies of the first three virtual orbitals obtained from HF, depicted in Figure 13. For this experiment we have used the cc-pVQZ basis, both directly and as a means to construct a DG basis. As a reference, we compare against the virtuals obtained using the aug-cc-pVQZ GTO basis in PySCF.
@@
$c$-pVDZ 73 -7.986163 0.001164 43 -7.985790 0.001547
$c$-pVTZ 150 -7.987134 0.000193 94 -7.986964 0.000373
$c$-pVQZ 180 -7.987292 0.000035 198 -7.987228 0.000109
@@
$c$-pVDZ 73 -7.918440 0.001156 43 -7.918077 0.001519
$c$-pVTZ 150 -7.919393 0.000203 94 -7.919217 0.000378
$c$-pVQZ 180 -7.919563 0.000033 198 -7.919493 0.000103
@@ 8.5 $\boldsymbol{\mathrm{H}_{2}\mathrm{O}}$
$c$-pVDZ 105 -76.035217 0.031826 51 -76.030384 0.036659
$c$-pVTZ 220 -76.058029 0.009014 114 -76.057214 0.009829
$c$-pVQZ 294 -76.065387 0.001656 255 -76.064823 0.002220
@@
$c$-pVDZ 105 -75.874622 0.038040 51 -75.867762 0.044899
$c$-pVTZ 220 -75.900554 0.012107 114 -75.910937 0.013143
$c$-pVQZ 294 -75.910075 0.002586 255 -75.909232 0.003429
@@ 8.6 $\boldsymbol{\mathrm{C}_{6}\mathrm{H}_{6}}$
For lower levels of GTO approximation, the induced DG basis is smaller than the ordinary GTO basis set, but the DG basis is actually smaller in the QZ case. Likewise, the DG basis sets for a given level of GTO approximation yield more accurate energies.
@@
$c$-pVDZ 476 -230.139441 0.062650 258 -230.135675 0.066417
$c$-pVTZ 966 -230.187508 0.014583 552 -230.184766 0.017326
$c$-pVQZ 1140 -230.199044 0.003047 1188 -230.198799 0.003292
@@ 9 Conclusion
Interestingly, we find that combining polynomial functions with GTOs in the provisional basis offers no significant accuracy gain over using GTOs alone in the HF and DFT settings. Nonetheless, polynomials offer a systematic path toward achieving a complete basis, which GTOs alone do not provide. An important future direction is to explore the question of systematic convergence for post-HF methods, where basis set incompleteness errors are more pronounced.